As at January 1, 2022, what is the value of P10,000 deposited on July 1, 2016, which accumulates interest at 16% compounded quarterly?
Answer #1
FV= PV*(1+r)^n
Where,
FV= Future Value
PV = Present Value
r = Interest rate =16%/4 =4%
n= periods in number =5.5 years *4 =22
= 1000*( 1+0.04)^22
=1000*2.36992
= P 2369.92